OpenStudy (anonymous):

Given f(x) = 3x-1/ 2 the quantity of 3x minus 1, divided by 2, solve for f-1(4). 1 please help medal 3 5 7

terenzreignz (terenzreignz):

Though of course, it can't hurt to try this one more time ^_^ \[\Large f(x) = \frac{3x - 1}{2}\]

terenzreignz (terenzreignz):

For now, it might be useful to find what's called the INVERSE of the function f. Do you know this term, inverse?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

btw i have to go in 3 min sry

OpenStudy (anonymous):

so trying to rush this

terenzreignz (terenzreignz):

I see. Then chop chop. There's a shortcut, in that case.

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

In plain English, you're actually being asked what value of x would yield a function value of 4. So, in that case, simply replace f(x) with 4, and then solve for x. \[\Large 4 = \frac{3x-1}{2}\] Hop to it.

OpenStudy (anonymous):

so plug in 4 for x

terenzreignz (terenzreignz):

No... plug 4 in for f(x), as I did, and then SOLVE for x.

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Go. The answer you get for x will be the answer to this question, it shouldn't take more than a minute. Hop to it :)

OpenStudy (anonymous):

im trying one sec

OpenStudy (anonymous):

not everyone as smart as you sadly :(

terenzreignz (terenzreignz):

Don't be silly... we're all smart~ just not equally wiling to admit it. Now... If you're really desperate and rushing, then just replace the x in \[\Large \frac{3x-1}{2}\] with 1,3,5, and 7 and see which one gives you a value of 4.

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

That is correct. Good job ^_^

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

have nice day